evaluate the expression
(-16)^1/4
hope i wrote this right

Answer by stanbon(75887)

evaluate the expression
(-16)^1/4
----------
Note: -1 = i^2 and 16=2^4
--------------------------------
So,(-16)^(1/4) = 2*sqrt(i)
===============================
Cheers,
Stan H.